1. What is the maximum frequency in 1018 Hz of photons produced in an CRT using a 20.65 kV accelerating potential, such as a color TV?

2. What is the maximum possible distance traveled in 14.2 μs by electrons ejected from a material by 77 nm photons, if they are bound to the material by 4.19 eV?

3. What is the wavelength of an electron moving at 3.27% of the speed of light? Give your answer in picometers.

4,What is the energy in 10-9 eV of a photon in a radio wave from an AM station that has a 1466 kHz broadcast frequency?

5. A physicist is watching a 10.4 kg orangutan at a zoo swing lazily in a tire at the end of a rope. He (the physicist) notices that each oscillation takes 4.15 s and hypothesizes that the energy is quantized. What is the difference in energy in 10-34 joules between allowed oscillator states?
Answer #1

1 )

using equation

f = E/ h

= qV/ h

= 1.6 X 10-19 X 20.65 X 103 / 6.67 X 10-34

f = 5.006 X 1018 Hz

2 )

KE = h C / \lambda - W

1/2 m v2 = h C / \lambda - W

= ( 6.67 X 10-34 X 3 X 108 / 1.6 X 10-19 ) / 77 - 4.19

= ( 6.67 X 10-34 X 3 X 108 / 1.6 X 10-19 ) / 77 - 4.19

= 1250/77 - 4.19

= 12.04 eV

v2 = 2 X 12.04 X 1.6 X 10-19 / 9.11 X 10-31

v = 2.0565 X 106 m/s

d = v X t

= 2.0565 X 106 X 14.2 X 10-6

d = 28.791 m

3 )

v = 3.27 %

= (3.27/100) X 3 X 108

v = 9810000 m/s

\lambda = h / m v

= 6.67 X 10-34 / 9.1 X 10-31 X 9810000

\lambda = 7.4716 X 10-11 m